OpenStudy (anonymous):

7.5/ 30% = please break it down

OpenStudy (anonymous):

1/400

OpenStudy (anonymous):

ok ill break it

OpenStudy (lalaly):

30% is 30/100=.3 so it becomes 7.5/.3

OpenStudy (anonymous):

my bad i got 25

OpenStudy (anonymous):

=7.5 / 0.3 = 25

OpenStudy (anonymous):

7/(30/100) 7/0.3

OpenStudy (anonymous):

which is also the sam as saying 750/30 or 75/3

OpenStudy (anonymous):

3/5 = (?)% * 5/8 please break it down